OpenStudy (asadkarim7):

Find the value of the coefficient of 1/x in the expansion of (2x − 1/x )^5 .

OpenStudy (asadkarim7):

i need explanation

OpenStudy (amistre64):

it looks like a binomial expansion, but the question itself seems a bit skewed

OpenStudy (amistre64):

\[{n\choose r} a^r\ b^{1-r}\] i think is the structure of it

OpenStudy (amistre64):

1/x is only alone in the last term of the expansion and I think I got the exponents backwards above \[{n\choose r} a^{n-r}\ b^{r}\] that looks better \[{5\choose 5} 2x^{5-5}\ (-1/x)^{5}\]

OpenStudy (amistre64):

(5 5) = 1 2x^0 = 1 (-1/x)^5 = -1 * 1/x

OpenStudy (mr.math):

The formula is \[(z+y)^n=\sum_{k=0}^n\left(\begin{matrix}n \\ k\end{matrix}\right) x^{n-k}y^k\] Now just substitute directly in the formula with \(n=5\) and \(k=1\) to find the coefficient of \(\frac{1}{x}\). Obviously \(y=-\frac{1}{x}\) and \(z=2x\).

OpenStudy (mr.math):

You need only to evaluate that term, you don't have to find all terms (i.e the term at k=1).

OpenStudy (amistre64):

{n \choose k} is simpler to type up

OpenStudy (amistre64):

\[{n \choose k}\]

OpenStudy (mr.math):

\[{n\choose k} \text{ Nice! Thanks :) }\]

OpenStudy (amistre64):

;)